OpenStudy (anonymous):

how does the expression (4^n)/(3^(n-1)) become 4(4/3)^n-1.?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}(4^n/3^{n-1})=\sum_{n=1}^{\infty}4(4/3)^{n-1}\] which simplifications were used